1. On January 1, 2020, Crane Company sold 17% bonds with a face value of $1900000. The bonds mature in five years, and interest is paid semiannually on June 30 and December 31. The bonds were sold for $2030500 to yield 15%. Using the effective-interest method of amortization, interest expense for 2020 is

a

$323000.

b

$304525.

c

$303885.
$285000.

2. On October 1, 2020 Sunland Company issued 4%, 10-year bonds with a face value of $5970000 at 104. Interest is paid on October 1 and April 1, with any premiums or discounts amortized on a straight-line basis.

Bond interest expense reported on the December 31, 2020 income statement of Sunland Company would be

a

$59700

b

$53730

c

$65670

d

$107460

3. A company offers a cash rebate of $2 on each $6 package of batteries sold during 2021. Historically, 10% of customers mail in the rebate form. During 2021, 5700000 packages of batteries are sold, and 177000 $2 rebates are mailed to customers. What is the rebate expense and liability, respectively, shown on the 2021 financial statements dated December 31?

a

$786000; $786000

b

$1140000; $786000

c

$354000; $786000

d

$1140000; $1140000
Answer #1

1. Interest expense for 2020 = $19,00,000*17/100

a) $ 3,23,000

2. Bond interest expense reported on the December 31, 2020 = $59,70,000*4%*3/12

a) $ 59,700

3. Computation of Rebate Expense and Liability

Rebate Expense = 57,00,000*10%*2=11,40,000

Liability = 11,40,000 - (1,77,000*2) = 7,86,000

b) $11,40,000; $7,86,000